Solve the proportion x-2/4=x+4/2
Answer by unlockmath(1688)

Hello,
The fastest way is cross multiply so it looks like this:
2(x-2)=4(x+4) Now distribute to look like this:
2x-4=4x+16 Add 4 to both sides and subtract 4x from both sides will give us:
-2x=20 Divide by -2
x=-10 Now let's put this into the original proportion and is should work out fine.
RJ Toftness
www.math-unlock.com